A sample of oxygen, O2, occupies 32.2 mL at 30 °C and 452 torr. What volume will it occupy at –70 °C and the same pressure?

1
Expert's answer
2021-12-20T21:16:01-0500

V1= 32.2ml

V2 =?

T1= 30°C +273 = 303K

T2 = -70°C +273 =203K


"\frac{V1}{T1}= \frac{V2}{T2}"


"=\frac{32.2ml}{303K}=\frac{V_2}{203K}"


"V_2=\frac{32.2ml × 203K}{303K}"


"V_2= 21.6ml"